a travel agency offers an organization an all-inclusive tour for $800 per person if nor more than 100 people take the tour. however the cost per person will be reduced to $5 for each person in excess of 100. how many people should take the tour in order for the agency to receive the largest gross revenue, and what is the largest gross revenue?

Answer by ikleyn(52814)

.
The condition says that in the domain n >= 100 the following formula works for the single ticket price P:

P = 800 - 5*(n-100)   dollars.


The revenue R is the product  R = n*P,   or

R = n*(800-5*(n-100)) = 800n - 5n^2 + 500n = 1300n - 5n^2 = 5n*(260-n).


The function R(x) = 5x*(260-x) is the parabola, and you need to find its maximum.


This parabola has the roots x = 0 and x = 260.


Therefore, its maximum is exactly mid-point between the roots, i.e. x = 130.


Thus, the optimal number of people in the tour is 130: this number gives the maximum to Revenue.


To calculate the value of this maximum, substitute x = n = 130 into the parabola equation

R(n) = 5n*(260-n) = 5*130*130   and calculate.


On finding the maximum of a quadratic function see the lessons
    - HOW TO complete the square to find the minimum/maximum of a quadratic function
    - Briefly on finding the minimum/maximum of a quadratic function
    - HOW TO complete the square to find the vertex of a parabola
    - Briefly on finding the vertex of a parabola
in this site.

On solving similar problems see the lesson
    - Using quadratic functions to solve problems on maximizing revenue/profit
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Finding minimum/maximum of quadratic functions".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.